Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 598     Entry time: 24 Mar 2025, 09:42
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 24.03.2025 
General Announcement:
- A. Galler informs that on Friday there is the LIMP meeting organized by Martin Doering. Please contact him in case of questions. Information will be transmitted to people who does not attend it.
- As PRC, A. Galler mentions the SXP laser shutter issue happened on Saturday (see SXP). 
- As DRC, David Lomidze informs it was a quiet week.

CTRL:
- Nothing to report.

DA:
- Not present.

EEE:
- Issue with CRL lenses in XTD2 with a stuck motor. A ZZ is needed tomorrow if possible to fix it.

VAC:
- Nothing to report.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Problems with GPFS system still in investigation.

XRO:
- Nothing to report.

LAS:
- Not present.

UNSYS:
- Nothing to report.

SPB:
- Nothing to report.

FXE:
- Data were collected; CRL got stuck.
- Pulse energy decreased ~20% every day and recovered in the morning; Also when SA2 changed number of pulses.
- uTCA crate crashed and a ticket is sent to EEE, together with DOC being called.

MID:
- Nothing to report.

HED:
- Not present.

SCS:
- Nothing to report.

SQS:
- Not present.

SXP:
- Saturday problem with Laser interlock related to AMPHOS shutter. A quick fix was found by PRC and Pouneh S.
- Responsibility for AMPHOS shutter will be discussed with LAS.
ELOG V3.1.4-7c3fd00